OpenStudy (anonymous):

A car traveling 88 km/h is 110 m behind a truck traveling 75 km/h. How long will it take the car to reach the truck?

OpenStudy (anonymous):

relative velocity is 80-75= 5km the time taken to cover distance 0.11km with velocity 5km/hr is the answer. . .

OpenStudy (anonymous):

i need the exact formula

OpenStudy (anonymous):

its 88 not 80

OpenStudy (anonymous):

Take the relative velocity 88 - 75 = 13Km/h change to SI unit 13 * 5/18 = 65/18 m/s Speed = Distance / time time = Distance / speed =110 * 18 / 65 = 30.46 s